OpenStudy (anonymous):

4-10y=22-6x in function form

OpenStudy (anonymous):

4-10y=22-6x Since 4 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 4 from both sides. -10y=-4-6x+22 Simplify the right-hand side of the equation. -10y=-6x+18 Divide each term in the equation by -10. -(10y)/(-10)=-(6x)/(-10)+(18)/(-10) Simplify the left-hand side of the equation by canceling the common factors. y=-(6x)/(-10)+(18)/(-10) Simplify the right-hand side of the equation by simplifying each term. y=(3x-9)/(5)

OpenStudy (anonymous):

medal please :)

jimthompson5910 (jim_thompson5910):

need one last step f(x) = (3x-9)/5 but you pretty much have it